# GuardSplat: Efficient and Robust Watermarking for 3D Gaussian Splatting

### v2

@@ C More Implementation Details
Table S1: Security analysis across various pruning ratios $K\%$. “Bottom $K$” denotes removing $K\%$ of the low-opacity Gaussians, while “random” denotes randomly removing $K\%$ of the Gaussians.
[…]
@@
D Additional Evaluation for Security
We conduct an additional experiments to evaluate the security of our GuardSplat in Table S1, investigating whether the malicious users can remove the watermarks from the model file by pruning the $K\%$ of Gaussians, where $K\in\{5,10,15,20,25\}$. “Bottom $K$” denotes pruning $K$ of low-opacity Gaussians, while “random” denotes randomly pruning $K$ of the Gaussians. As demonstrated, our GuardSplat still achieves a bit accuracy of 98.74% when 25% of the low-opacity Gaussians are removed, indicating that simply removing low-opacity Gaussians does not effectively attack our method. Though randomly removing the Gaussians can lead to a significant decline in bit accuracy, it also greatly affects the reconstruction quality (i.e., PSNR, SSIM, and LPIPS), resulting in low-fidelity rendering. This experimental result demonstrates that the malicious cannot directly remove the watermarks from the model file, verifying the security of our GuardSplat.
Figure S3: Visual comparisons between various watermarking embedding strategies and our SH-aware module. Heatmaps at the bottom show the differences (×$\times 10$) between the watermarked and Groundtruth. Red text indicates the best performance.
@@ E More Results
E.1 Additional Visual Comparisons
Various Watermark Embedding Strategies. In the main paper, we explore the performance under various watermark embedding strategies with $N_{L}=32$ bits (see quantitative results in Table 3). For better comparisons, we further visualize the results of various watermark embedding strategies in Figure S3. As shown, the proposed SH-aware module achieves superior bit accuracy and reconstruction quality to the competitors.
Various Loss Combinations. In the main paper, we quantitatively compare the performance across various loss combinations in Table 4. We also conduct a visual comparison of these ablation variants in Figure S4. As shown, “$\mathcal{L}_{\text{\text{recon}}}+\mathcal{L}_{\text{\text{msg}}}+\mathcal{L}_{\text{\text{off}}}$” achieves the best performance in bit accuracy and reconstruction quality.
Table S2: Quantitative results of our GuardSplat on Blender [27] and LLFF [26] datasets w.r.t $N_{L}\in\{64,72\}$ bits.
@@ E.1 Additional Visual Comparisons
E.2 Quantitative Results on Larger-Capacity Messages.

### v5

@@ D.2 Various Loss Combinations
Table S1: Security analysis across various pruning ratios $\mathtt{K}\%$. Bottom K denotes removing $\mathtt{K}\%$ of the low-opacity Gaussians, while Random denotes randomly removing $\mathtt{K}\%$ of the Gaussians.
% Bottom $K$ Random
@@
Figure S3: Visual comparisons between various message embedding strategies and our SH-aware module. Heatmaps at the bottom show the differences (×$\times 10$) between the watermarked and Groundtruth. Red text indicates the best performance.
@@ E More Results
E.1 Quantitative Results on Larger-Capacity Messages
E.2 Bit Accuracy across Various Rendering Situations
We explore the extraction accuracy of learned SH offsets across the following situations: 1) SH Noise; 2) Light Conditions; 3) Occlusions; and 4) Viewing Angles. Specifically, to simulate different lighting conditions, we first train a 3DGS asset of “Lego” from the TensoIR [16] dataset in "RGBA" mode. We then freeze all Gaussian attributes while optimizing the SH features to adapt to various illumination scenarios, such as “light”, “sunset”, and “city”. We train only the SH offsets in “RGBA” mode and add them to the SH features of other lighting modes for evaluation. As shown in Figure S5, GuardSplat achieves good robustness against SH noise (a) and light conditions (b) by adding noise to SH features in training. Since the occluded areas can be removed by segmentation models (e.g., Segment Anything Model [21], and Grounding DINO [25]), we train GuardSplat to extract messages from randomly masked views ($\leq 20\%$). It improves the robustness of our GuardSplat against various occlusions (c). GuardSplat is inherently robust to various viewing angles (d) since it is designed for 3D.
Table S2: Quantitative results of our GuardSplat on Blender [32] and LLFF [31] datasets with $N_{L}\in\{64,72\}$ bits.
@@ E.2 Bit Accuracy across Various Rendering Situations
Figure S5: Bit accuracy across various rendering parameters.
Figure S6: Zoomed-in rendering results between the original 3DGS and our GuardSplat.
E.3 Zoomed-in Rendering Results
Since SH features produce highly realistic shading and shadowing, altering them may reduce fidelity, especially in the specular areas. To clearly show how the SH offsets are changing the rendering results, we conduct a visual comparison of zoomed-in rendering results between the original 3DGS and our GuardSplat of “ball” on the Shiny [54] dataset in Figure S6. As shown, GuardSplat can preserve the original metallic luster of assets.